OpenStudy (anonymous):

sqrt{x^6y^10} sqrt{x^6y^10} @Mathematics

OpenStudy (anonymous):

\[\sqrt{x ^{6}}y ^{10}\]

OpenStudy (anonymous):

ok i didnt get that answer. how did you work it?

OpenStudy (anonymous):

\[\sqrt{x ^{6}}=(x ^{6})^{1/2} = x ^{6/2} = x ^{3}\]

OpenStudy (anonymous):

Same method for y also.

OpenStudy (anonymous):

how do you know to use 1/2? thats where i messed up

OpenStudy (anonymous):

For square we use ^2, for sqrt we use ^1/2

OpenStudy (anonymous):

\[\sqrt{16} = 16^{1/2}\]

OpenStudy (anonymous):

thats confusing lol

OpenStudy (anonymous):

\[\sqrt{2}*\sqrt{2} = 2\] \[2^{1/2}*2^{1/2}=2^{1/2+1/2}=2^{1} = 2\]

OpenStudy (anonymous):

Anymore confusion?

OpenStudy (anonymous):

no i think im good. just trying to check my work and its just hard because of trying to type the problem in the room so people can understand what im asking lol